OpenStudy (anonymous):

write a recursive formula for this sequence: 1,4,10,22

OpenStudy (jdoe0001):

I was thinking.... this sorta works \(\large \begin{array}{ccccc} 1&2&3&4&5 \\ \hline\\ 1&4&10&22&\square \\ \hline\\ &2(2-1)+2&2(3-1)+2&2(4-1)+2&2(n-1)+2\\ \end{array}\)

OpenStudy (jdoe0001):

hmm.. that ... one sec

OpenStudy (anonymous):

so that would be my recrusive formula?

OpenStudy (jdoe0001):

\( \begin{array}{ccccc} 1&2&3&4&5 \\ \hline\\ 1&4&10&22&\square \\ \hline\\ &2(2^{th}-1)+2&2(3^{th}-1)+2&2(4^{th}-1)+2&2(n^{th}-1)+2 \end{array}\)

OpenStudy (anonymous):

what would be my explicit formula?

OpenStudy (jdoe0001):

is more like a arithmetic series

OpenStudy (jdoe0001):

well, is 2 times current term value minus 1 plus 2 so say.. if you're in the 2nd term 2nd term - 1, gives the 1st term, and thus its value 1st term value is 1 2nd will be 2 ( 1-1) +2 = 2 3rd will be 2 ( 2-1) + 2 = 4 4th will be 2( 4-1) + 2 = .....

OpenStudy (jdoe0001):

hmm.... well one sec

OpenStudy (jdoe0001):

well, is 2 times current term value minus 1 plus 2 so say.. if you're in the 2nd term 2nd term - 1, gives the 1st term, and thus its value 1st term value is 1 2nd term will be 2(2nd - 1)+2 = 2 ( 1)+2 = 4 3rd term will be 2(3rd - 1) + 2 = 2( 4) + 2 = 10 4th term will be 2(4th -1)+2 = 2(10)+2 = 22

OpenStudy (jdoe0001):

(n-1) meaning..... the current "n term" -1 so that'd give the previous term times 2 +2

OpenStudy (anonymous):

no it's actually set up like this 0,1 1,4 2,10 3,22

OpenStudy (anonymous):

its not 1,1

OpenStudy (jdoe0001):

well, it'd still apply \(\large \begin{array}{ccccc} 0&1&2&3&4 \\ \hline\\ 1&4&10&22&\square \\ \hline\\ &2(1^{st}-1)+2&2(2^{nd}-1)+2&2(3^{rd}-1)+2&2(n^{th}-1)+2 \end{array}\)

OpenStudy (jdoe0001):

the 1st term will then be 2( 1st -1) 2(1) + 2 = 4

OpenStudy (anonymous):

1-1=0

OpenStudy (jdoe0001):

the (1-1) term, that is, the 0 term value is 1

OpenStudy (anonymous):

okay so what's the recursive formula? im a little confused

OpenStudy (jdoe0001):

the "n" stands for the CURRENT term then you subtract 1 from it, gives you the PREVIOUS term so (1-1) should give you the \(\bf 0^{th}\) term and thus its value, of 1

OpenStudy (jdoe0001):

2(2nd - 1) will give you the 2 ( 2-1) or the 1st term, 2(1st term) whose value is 4 2(4)+2 = 10

OpenStudy (anonymous):

so the formula is 2(nth-1)+2?

OpenStudy (jdoe0001):

skipping the 1st term, I'd say yes, that'd the SERIE

OpenStudy (anonymous):

but I can't skip the first term because it's apart of the series

OpenStudy (anonymous):

how would that work out?

OpenStudy (jdoe0001):

was thinking about that..... I figure..... the 1st one is included hmm

OpenStudy (jdoe0001):

hmmm I don't see the inclusion of the 1st term .... simply enough.....

OpenStudy (anonymous):

ok thanks